OpenStudy (anonymous):

A house was valued at $125,000 in the year 1990. The value appreciated to $170,000 by the year 2008. A) What was the annual growth rate between 1990 and 2008? = Round the growth rate to 4 decimal places. B) What is the correct answer to part A written in percentage form? = %. C) Assume that the house value continues to grow by the same percentage. What will the value equal in the year 2012 ? value = $ Round to the nearest thousand dollars.

OpenStudy (anonymous):

sorry guys these are the ones im not sure about at all:(

OpenStudy (anonymous):

C = S * (1 + P)^Y where C = current value, S = Starting Value, P = Growth Rate, and Y = years Therefore, \[P = (C/s)^{1/Y} -1 = .01722922 = .01723\] To put as a percentage, multiply by 100: 1.723% Use the formula again, but you don't know C \[C = s*(1+P)^Y = $182022.135\]

OpenStudy (anonymous):

sorry i dont know how you got the .01722922=.01723? i thought you put (125000/170000)^1-1 would give you the answer but it doesnt:(

OpenStudy (anonymous):

It's 170000/125000 :D